# Efficient stripe artefact removal by a variational method: application to light-sheet microscopy, FIB-SEM and remote sensing images

### v1

@@ 4 Conclusion
In this paper, we focused on the application of different stripe removal methods on LSFM images and a subsequent evaluation of results. We explained Fourier filtering and variational methods using selected methods. These incorporated well-founded extensions of existing propositions for increased quality of stripe removal. Afterwards, we used synthetically generated LSFM data and real images to evaluate and compare the performance of stripe removal by using visual inspection. The observations are confirmed using quality metrics. This revealed that our proposed objective function (6), later referred to as GSR, produces consistent results with better stripe removal than prior propositions.
For Fourier filtering, we used a modified version of MDSR. The method applies a structural decomposition by shape and scale called NSCT. Afterwards, components with directions close to stripe direction are reduced by applying appropriate filtering in the Fourier domain where image and stripe information are more locally separated. A transformation back towards the image domain yields the result. On the other hand, for variational methods we discussed the construction of the objective function of GSR in detail. Individual terms are formulated by encoding properties of stripes and structures. These are then combined into a weighted functional and optimized by an minimization algorithm.
During our evaluation on real and synthetic data we found out that the GSR introduced in this paper produces consistently good results across all examples. It offers the best removal of stripe artifacts for all variations of observed stripes. This includes stripes of varying lengths, widths, intensities and slightly deviated directions. Other methods such as the previously published MDSR and VSNR produce a lower reduction of artifacts and show inconsistencies when faced by different image appearances and structures.
@@
The usage of synthetic LSFM data which offer a stripe-free reference image is a novel supplement of real data for comparing quality of stripe removal. It provides valuable insights into the capabilities and limitations of stripe removal methods. Furthermore, it allows for the use of standard image quality metrics such as PSNR and MS-SSIM which further assist in a precise assessment.
The curtaining metric proved to be a valuable addition to visual inspection as indicator for stripe corruption. Despite its shortcomings curtaining values reaffirm our visual inspection and successfully convey a reduction in stripe artifacts in almost all cases. Its major advantages are that values are calculated directly on the image without a need of reference and that it measures corruptions by stripes. However, visual inspection remains indispensable in evaluating results.
Appendix A LSFM Simulation
The creation of synthetic LSFM data such as shown in Figure 15 requires the simulation of light propagation through media to capture scattering and absorption effects correctly. The python package biobeam [18] provides the necessary functions for calculating light propagation through a volume of media. The geometric properties of the volume are specified through a distribution of refractive indices (rid) and a fluorescence distribution (fld). The rid is denoted by $\mathbf{n}\in\mathbb{C}^{n_{x}\times n_{y}\times n_{z}}$ and contains complex valued refractive indices in each pixel. The real part is the classical refractive index and describes scattering behaviour while the imaginary part is the absorption coefficient which describes effects such as light attenuation. For generating reference images without light interaction we set $\mathbf{n}\equiv n_{0}$ to be constant. A reasonable choice is $n_{0}=1.33$ which is approximately the refractive index of water which is the main constituent of biological cells.
The fld is denoted by $f\in\mathbb{R}_{\geq 0}^{n_{x}\times n_{y}\times n_{z}}$. It contains multiplicative non-negative factors that signify the proportion of re-emitted light by fluorescence when subjected to illumination. The fld is a simplified representation of reality. Typically, biological samples are labeled for LSFM using a fluorescent marker, which are specialized proteins that emit light when illuminated at certain wavelengths. The intensity of the emitted fluorescence is linked to the concentration of marker protein within the sample.
Figure 15: Simulated images (a-c) and references (d-f) for different simulation examples. Cell cluster (a,d), dispersed cells (b,e) and embryo (c,f).
Besides geometrical properties of a synthetic sample, we have to choose parameters of the optical setup. We chose the wavelength of illumination light $\lambda=500\mathrm{nm}$, the numerical apertures of illumination and detection optics $\mathrm{NA}_{\mathrm{ill}}=0.1$, $\mathrm{NA}_{\mathrm{det}}=0.4$ and the voxel size $v=[0.4,0.4,0.4]\mu\mathrm{m}$. The corresponding values were fixed for all synthetic image data in this paper. Furthermore, we chose the illumination direction along the $y$-axis and the detection direction along the $z$-axis.
With a given setup of geometry and imaging, simulations of synthetic LSFM images can be obtained. The following steps model the real imaging process:
  • (i)
    Illumination: From $\mathrm{NA}_{\mathrm{ill}}$ and $\lambda$ a complex-valued $x$-$z$ cross-section of a cylindrical light sheet is calculated. Its height $\bar{z}$ corresponds to the illuminated slice in the sample and can be chosen arbitrarily. With wave-optical propagation as described in [7], the cross-section is propagated through the sample. This yields the illuminated volume $I_{\mathrm{ill}}$.
  • (ii)
    Fluorescence: Afterwards, the amount of light re-emitted through fluorescence is obtained via
    $I_{\mathrm{fl}}=I_{\mathrm{ill}}\odot f$ (11)
    with element-wise multiplication $\odot$. It represents the ’fluorescence response’ to the illumination.
  • (iii)
    Acquisition: From $\mathrm{NA}_{\mathrm{det}}$, a detection point-spread function (PSF) $h_{\mathrm{det}}$ is calculated. It is invariant under rotations around the detection axis. A 2D image of the illuminated slice is obtained by convolution with this PSF and extraction of the corresponding slice
    $I_{\mathrm{det}}=(I_{\mathrm{fl}}\ast h_{\mathrm{det}})_{z=\bar{z}}.$ (12)
  • (iv)
    Repeating steps (i)-(iii) for light sheet heights $\bar{z}\in\{1,\dotsc,n_{z}\}$, slices are combined into a 3D volume image.
Figure 16: Visualization of the simulation procedure. $|\mathrm{rid}|$ (a), fld (b), $I_{\mathrm{ill}}$ (c), $I_{\mathrm{fl}}$ (d), $I_{\mathrm{det}}$ (e), $h_{\mathrm{det}}$ (f). (a)-(e) are $y$-$x$-cross-sections of the illuminated slice extracted from the volume. (f) is a $z$-$x$-cross-section of $h_{\mathrm{det}}$ with 4x magnification compared to (a)-(e).
The essential parts of the simulation procedure are visualized in Figure 16. This process is a simplified model of reality. For example, we use idealized versions of the optical systems. Therefore, the cylindrical light sheet is of constant thickness along the $x$-direction and the detection PSF remains constant for all pixels. In real systems, these approximations are reasonable for the area of interest but do not hold true further away from it.
The synthetic data shown in Figure 15 depicts some examples of simulated LSFM images with an appearance similar to real data. All samples were created by using non-overlapping spheres as main foreground. In addition, small variations inside of the spheres were created by adding fine scale 3D Perlin noise. For some structural background a coarse scale 3D Perlin noise was added, except for the dispersed cells image. Values of the rid for different structures were chosen based on reference values [40]. On the other hand, values for the fld were chosen to produce visually appealing images which are in accordance with real LSFM images.
Appendix B Parameter Selection for GSR
In this section of the appendix, we want to give insight into the parameters of the GSR algorithm. More precisely, we want to give advice on how parameters can be selected and tuned for an image. Therefore, we discuss the behaviour of varying $\mu_{1},\mu_{2}$ and propose two sets of parameters for producing an initial result depending on the problem at hand. Fine-tuning towards a more desirable result may still be necessary and can then be done according to our following observations.
Figure 17 displays the effects of varying $\mu_{1}$ and $\mu_{2}$ for a selection of parameters with reasonable outcomes. We notice that the rules of thumb established in Subsection 2.2 can be observed here. More precisely, an increase in $\mu_{1}$ corresponds to a stronger reduction in stripes which can lead to smoothing and smearing artifacts as most dominantly shown in (h). On the other hand, increasing $\mu_{2}$ reduces the amount of stripe reduction and counteracts these kinds of artifacts. However, scaling of both $\mu_{1}$ and $\mu_{2}$ by an equal factor, e.g., compare (b) and (i), produces different results. In our experience, scaling with a factor $>1$ increases stripe removal with minimal effect on introducing artifacts. Nonetheless, this is only valid in a reasonable range of parameters which we propose to be approximately $\mu_{1}\in[0.15,0.5]$ and $\mu_{2}\in[0.0033,0.017]$. Finding ’optimal’ results requires some initial guess followed by manual fine-tuning. In all considered cases choosing either $(\mu_{1},\mu_{2})=(1/3,1/300)$ or $(\mu_{1},\mu_{2})=(1/2,1/30)$ produced sufficiently good results. The larger parameters usually perform better when stripe artifacts appear in a more difficult setting, e.g., when they are short, wide or slightly oblique. In general, $(\mu_{1},\mu_{2})=(1/3,1/300)$ yielded superior results when stripes were thin, long or had smoothed out edges. In our case, fine-tuning was rarely necessary and can be performed according to the previous observations and the rules of thumb in Subsection 2.2. If more stripe removal is wanted, scaling both parameters by an equal factor is a good start. Alternatively, increasing only $\mu_{1}$ can achieve the same effect. If image structures are too strongly affected, choosing a larger $\mu_{2}$ typically resolves the issue.
Figure 17: Results of GSR(2D) on the mouse intestine cuboid with different parameter settings. Image (a), $\mu=(\frac{1}{6},\frac{1}{300})$ (b), $\mu=(\frac{1}{6},\frac{1}{100})$ (c), $\mu=(\frac{1}{6},\frac{1}{60})$ (d), $\mu=(\frac{1}{3},\frac{1}{300})$ (e), $\mu=(\frac{1}{3},\frac{1}{100})$ (f), $\mu=(\frac{1}{3},\frac{1}{60})$ (g), $\mu=(\frac{1}{2},\frac{1}{300})$ (h), $\mu=(\frac{1}{2},\frac{1}{100})$ (i), $\mu=(\frac{1}{2},\frac{1}{60})$ (j).
@@ Acknowledgments
We thank Martin Weigert for guidance on installing biobeam and modelling LSFM imaging with it and Michael Engstler for providing the FIB-SEM image data. Furthermore, we want to thank Diego Roldán for providing and explaining his proposed quality metrics and Jannik Reiser for his help with processing results.
@@
The data and methods presented in this paper are available as supplementary material. This includes MATLAB code for all used methods and scripts for processing images. Furthermore, a Jupyter-notebook for generating synthetic LSFM data and a txt-file for creating the Python environment on Windows 11 are available.

### v2

@@ 4 Conclusion
In this paper, we evaluated several stripe removal methods on LSFM, FIB-SEM and remote sensing images. We explained Fourier filtering and variational methods using selected examples. Extensions of existing propositions increased the quality of stripe removal. We used synthetically generated LSFM data and real images to evaluate and compare the performance of stripe removal by visual inspection. The observations are confirmed by quality metrics. Our proposed objective function (6), later referred to as GSR, produced consistent results with better stripe removal than prior propositions. This includes stripes of varying lengths, widths, intensities and slightly deviated directions. Other methods such as the previously published MDSR and VSNR produce a lower reduction of artifacts and show inconsistencies when facing different image appearances and structures.
@@
The usage of synthetic LSFM data which offer a stripe-free reference image is a novel supplement of real data for comparing quality of stripe removal. It provides valuable insights into the capabilities and limitations of stripe removal method, e.g., inconsistent shape reconstruction, insufficient stripe removal or the introduction of smearing artifacts. Furthermore, it allows for the use of standard image quality metrics such as PSNR and MS-SSIM which further assist in a precise assessment. The recently proposed curtaining metric proved to be an valuable indicator of stripe corruption. Its major advantage is that it is calculated directly on the image and does not require a reference. However, it does not replace the need for visual inspection.
Supplemental document
See Supplement 1 for supporting content.
@@ Acknowledgments
We thank Martin Weigert for guidance on installing biobeam and its application for modelling LSFM imaging as well as Michael Engstler for providing the FIB-SEM image data. Furthermore, we want to thank Diego Roldán for providing and explaining his curtaining metric and Jannik Reiser for his help with data processing.
@@
The methods presented in this paper will be made available. This includes MATLAB code for all used methods and scripts for processing images. Furthermore, a Jupyter-notebook for generating synthetic LSFM data and a txt-file for creating the Python environment on Windows 11 are provided.